# On the isomorphism problem for ultraproducts of $\mathrm{C}^{*}$-algebras in continuous model theory

### v1

@@ 1 Introduction
In model theory, the isomorphism of ultraproducts of two structures is known to be closely related to a property called elementary equivalence. One of the earliest results in this direction was given by Keisler [Keisler:1961], which states that, under the assumption of the continuum hypothesis, for any $\mathcal{L}$-theory $\mathrm{T}$ in a countable language $\mathcal{L}$ and any two models $M$ and $N$ of $\mathrm{T}$ of size $\leq{\mathfrak{c}}$, $M$ and $N$ are elementarily equivalent if and only if they have isomorphic ultrapowers with respect to an ultrafilter on $\omega$. Furthermore, this result was generalized by Shelah [Shelah:1971], who weakened the assumption on the ultrafilter from one on $\omega$ to an ultrafilter on an arbitrary index set, and showed that the theorem holds even in ZFC. This is known as the Keisler-Shelah theorem. Moreover, it is also known to hold in the framework of continuous model theory - see [MTFMS, Theorem 5.7]. Thus, the following question naturally arises.
@@
Assume that the continuum hypothesis fails. For which $\mathcal{L}$-theories $\mathrm{T}$ in a countable language $\mathcal{L}$, do there exist models $M$ and $N$ of $\mathrm{T}$ of size $\leq{\mathfrak{c}}$ (or $\aleph_{0}$) such that $M$ and $N$ are elementarily equivalent, but $M^{\mathcal{U}}$ and $N^{\mathcal{V}}$ are not isomorphic for all nonprincipal ultrafilters $\mathcal{U},\mathcal{V}$ on $\omega$?
@@
Assume that the continuum hypothesis fails. There exists a $\mathrm{C}^{\ast}$}-algebra $M$ satisfying the following. We can find models $A$ and $B$ of the theory of $M$ with density characters $\leq{\mathfrak{c}}$ such that $A^{\mathcal{U}}$ and $B^{\mathcal{V}}$ are not isomorphic for all nonprincipal ultrafilters $\mathcal{U},\mathcal{V}$ on $\omega$.
@@ 2.1 Languages and Structures
A $\mathrm{C}^{\ast}$}-algebra is a complex Banach algebra $A$ equipped with an involution $\ast\colon A\to A$ (i.e., a map satisfying $(\lambda a+b)^{\ast}=\bar{\lambda}a^{\ast}+b^{\ast}$, $(a^{\ast})^{\ast}=a$, and $(ab)^{\ast}=b^{\ast}a^{\ast}$ for all $a,b\in A$ and $\lambda\in\mathbb{C}$) such that the $\mathrm{C}^{\ast}$}-identity $\left\lVert{a^{\ast}a}\right\rVert=\left\lVert{a}\right\rVert^{2}$ holds for all $x\in A$. $\mathrm{C}^{\ast}$}-algebras are not bounded, so they cannot be regarded as metric structures as they are. To address this, we introduce the following language and structure so that they can be treated as metric structures.
@@ 2.4 The Downward Löwenheim-Skolem-Tarski Theorem and Elementary chains
If $(I,<)$ is an ordinal $\alpha$ with $\mathrm{cf}(\alpha)\neq\omega$ (in particular, when $\alpha$ is a regular uncountable cardinal such as $\omega_{1}$ and $\omega_{2}$), then the union of an elementary chain $(M_{\beta})_{\beta<\alpha}$ is automatically complete, since any Cauchy sequences are contained in some $M_{\beta}$.
@@ 3.4 Definable Predicates and Definable Sets
4 Some Results under Continuum Hypothesis
It is well known that countable ultrapowers exhibit good behavior under the assumption of the Continuum Hypothesis (CH). In this section, we present several results in this direction. For simplicity, we assume throughout this section that the language $\mathcal{L}$ is separable with respect to the theory $\mathrm{T}$.
@@ 4 Some Results under Continuum Hypothesis
Suppose that $M$ and $N$ are saturated models of $\mathrm{T}$ that are elementarily equivalent and have the same uncountable density character $\kappa$. Then $M$ and $N$ are isomorphic.
@@
Let $M$ and $N$ be non-compact models of $\mathrm{T}$ that are elementarily equivalent and satisfy $\chi(M),\chi(N)\leq{\mathfrak{c}}$. Then, for all $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$, the ultrapowers $M^{\mathcal{U}}$ and $N^{\mathcal{V}}$ are isomorphic.
@@
By section 3.3, $M^{\mathcal{U}}$ is ${\mathfrak{c}}=\aleph_{1}$-saturated under CH holds. Thus, by virtue of section 4, it suffices to show that the density character of $M^{\mathcal{U}}$ is equal to $\mathfrak{c}$}. By the construction of the ultraproduct, we have $\chi(M^{\mathcal{U}})\leq\chi(M)^{\aleph_{0}}\leq{\mathfrak{c}}^{\aleph_{0}}={\mathfrak{c}}$. On the other hand, by section 3.1, we can regard $M$ as a substructure of $M^{\mathcal{U}}$, and hence as a closed subset of $M^{\mathcal{U}}$. Thus, if $M$ is non-compact, then $M^{\mathcal{U}}$ is also non-compact. Therefore, by section 3.3, we have that $\chi(M^{\mathcal{U}})\geq\aleph_{1}={\mathfrak{c}}$. Hence, we obtain $\chi(M^{\mathcal{U}})={\mathfrak{c}}$. ∎
@@
Let $M$ and $N$ be saturated models of $\mathrm{T}$. Suppose that $A\preceq M$ with $\chi(A)<\chi(M)$ and $\Phi\colon A\to N$ is an elementary embedding. Then, there exist two distinct partial elementary embeddings $\Phi_{0},\Phi_{1}\colon M\to N$ which both extend $\Phi$.
@@
Let $M$ and $N$ be saturated models of $\mathrm{T}$. Suppose $A\preceq M$ with $\chi(A)<\chi(M)$, and let $\Phi\colon A\to N$ be an elementary embedding. Then, there exist at least $2^{\chi(M)}$ isomorphisms from $M$ to $N$ that extend $\Phi$.
@@
Assume CH holds. Suppose $M$ and $N$ are non-compact separable models of $\mathrm{T}$, and let $\Phi\colon M\to N$ be an isomorphism. Then, for any $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$, there exist $2^{\mathfrak{c}}$ isomorphisms from $M^{\mathcal{U}}$ to $N^{\mathcal{V}}$ that extend $\Phi$.
@@
Assume CH holds. Let $A,B$ be infinite-dimensional separable unital $\mathrm{C}^{\ast}$}-algebras and $\Phi\colon A\to B$ an isomorphism. Then, for any $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$, there exist at least $2^{\mathfrak{c}}$ isomorphisms from $A^{\prime}\cap A^{\mathcal{U}}$ to $B^{\prime}\cap B^{\mathcal{V}}$.
@@ 5 Main Result
For a $\mathrm{C}^{\ast}$}-algebra AA, $\mathcal{P}(A)$ denotes the set of all projections in $A$. For $\in\mathcal{P}(A), pp$ is said to be minimal if, for any $q\in\mathcal{P}(A)$ with $0\leq q\leq p$ one has $q=0$ or $q=p$.
We let $\mathcal{C}$ denote the class of all unital $\mathrm{C}^{\ast}$}-algebras AA such that $\mathcal{P}(A)$ is commutative with respect to the product in $A$ and $A$ has no minimal projections.
@@
The class $\mathcal{C}$ is axiomatizable.
@@
By section 3.4, $\mathcal{P}$ is definable. It follows from section 3.4 that
$\sup_{p\in\mathcal{P}}\sup_{q\in\mathcal{P}}\left\lVert{pq-qp}\right\rVert$ (5.1.1)
is a definable predicate relative to $\mathrm{T}_{\mathrm{C}^{\ast}}$. Thus, the commutativity of projections can be described by the $\mathcal{L}_{{\mathrm{C}^{\ast}}}$-formulas.
Suppose that $A$ is a unital $\mathrm{C}^{\ast}$}-algebra with $\mathcal{P}(A)$ commutative. For $p,q\in\mathcal{P}(A)$, the condition $q\leq p$ is equivalent to $pq=q$, and thus $\left\lVert{pq-q}\right\rVert=0$. The product of projections is again a projection since $\mathcal{P}(A)$ is commutative. Hence, if $\left\lVert{pq-q}\right\rVert\neq 0$, then $\left\lVert{pq-q}\right\rVert=\left\lVert{(1-p)q}\right\rVert=1$, for $p,q\in\mathcal{P}(A)$. Similarly, for all $p,q\in\mathcal{P}(A)$ with $q\leq p$, if $\left\lVert{p-q}\right\rVert\neq 0$, then $\left\lVert{p-q}\right\rVert=\left\lVert{p-pq}\right\rVert=1$. Therefore, $A\in\mathcal{C}$ if and only if the interpretation of the following definable predicate
$\sup_{p\in\mathcal{P}}\sup_{q\in\mathcal{P}}\min\{1-\left\lVert{pq-q}\right\rVert,\left\lVert{q}\right\rVert,\left\lVert{p-q}\right\rVert\}$ (5.1.2)
is equal to 0. Hence $\mathcal{C}$ is axiomatizable. ∎
@@
Let $X$ be the Cantor space $2^{\omega}$. The $\mathrm{C}^{\ast}$}-algebra $A=C(X)$ is an abelian unital $\mathrm{C}^{\ast}$}-algebra with no minimal projections. More generally, for any abelian unital $\mathrm{C}^{\ast}$}-algebra $A=C(X)$, where $X$ is a compact Hausdorff space, $A$ has no minimal projections if and only if the Boolean algebra $\operatorname{clop}(X)$ consisting of all clopen subsets of $X$ is atomless.

### v2

@@ 1 Introduction
In model theory, the isomorphism of ultraproducts of two structures is known to be closely related to a property called elementary equivalence. One of the earliest results in this direction was given by Keisler [Keisler:1961], which states that, under the assumption of the Continuum Hypothesis, for any $\mathcal{L}$-theory $\mathrm{T}$ in a countable language $\mathcal{L}$ and any two models $M$ and $N$ of $\mathrm{T}$ of size $\leq{\mathfrak{c}}$, $M$ and $N$ are elementarily equivalent if and only if they have isomorphic ultrapowers with respect to an ultrafilter on $\omega$. Furthermore, Shelah [Shelah:1971] (see also [Shelah:1992]) generalized this result by weakening the assumption on the ultrafilter from $\omega$ to enough large cardinals, thereby showing that the theorem holds in ZFC. This is known as the Keisler-Shelah theorem. Moreover, it is also known to hold in the framework of continuous model theory - see [MTFMS, Theorem 5.7]. Thus, the following question naturally arises.
@@
Assume the Continuum Hypothesis fails. For which $\mathcal{L}$-theories $\mathrm{T}$ in a countable language $\mathcal{L}$, do there exist models $M$ and $N$ of $\mathrm{T}$ of size $\leq{\mathfrak{c}}$ (or $\aleph_{0}$) such that $M$ and $N$ are elementarily equivalent, but $M^{\mathcal{U}}$ and $N^{\mathcal{V}}$ are not isomorphic for all nonprincipal ultrafilters $\mathcal{U},\mathcal{V}$ on $\omega$?
@@
Assume the Continuum Hypothesis fails. Then, there exists a complete extension $\mathrm{T}$ of the theory of C*-algebras such that there are models $A$ and $B$ of $\mathrm{T}$ with density characters at most $\mathfrak{c}$} for which $A^{\mathcal{U}}$ and $B^{\mathcal{V}}$ are not isomorphic for any nonprincipal ultrafilters $\mathcal{U},\mathcal{V}$ on $\omega$.
@@ 2.1 Languages and Structures
A $\mathrm{C}^{\ast}$}-algebra is a complex Banach algebra $A$ equipped with an involution $\ast\colon A\to A$ (i.e., a map satisfying $(\lambda a+b)^{\ast}=\bar{\lambda}a^{\ast}+b^{\ast}$, $(a^{\ast})^{\ast}=a$, and $(ab)^{\ast}=b^{\ast}a^{\ast}$ for all $a,b\in A$ and $\lambda\in\mathbb{C}$) such that the $\mathrm{C}^{\ast}$}-identity $\left\lVert{a^{\ast}a}\right\rVert=\left\lVert{a}\right\rVert^{2}$ holds for all $a\in A$. $\mathrm{C}^{\ast}$}-algebras are not bounded, so they cannot be regarded as metric structures as they are. To address this, we introduce the following language and structure so that they can be treated as metric structures.
@@ 2.4 The Downward Löwenheim-Skolem-Tarski Theorem and Elementary chains
If $(I,<)$ is an ordinal $\alpha$ with $\operatorname{cf}(\alpha)\neq\omega$ (in particular, when $\alpha$ is a regular uncountable cardinal such as $\omega_{1}$ and $\omega_{2}$), then the union of an elementary chain $(M_{\beta})_{\beta<\alpha}$ is automatically complete, since any Cauchy sequences are contained in some $M_{\beta}$.
@@ 3.4 Definable Predicates and Definable Sets
4 Some Results under the Continuum Hypothesis
It is well known that countable ultrapowers exhibit good behavior under the assumption of the Continuum Hypothesis. In this section, we present several results in this direction. For simplicity, we assume throughout this section that the language $\mathcal{L}$ is separable with respect to the complete theory $\mathrm{T}$.
@@ 4 Some Results under the Continuum Hypothesis
Suppose that $M$ and $N$ are saturated models of $\mathrm{T}$ that have the same uncountable density character $\kappa$. Then $M$ and $N$ are isomorphic.
@@
Assume that the Continuum Hypothesis holds. Let $M$ and $N$ be non-compact models of $\mathrm{T}$ that are elementarily equivalent and satisfy $\chi(M),\chi(N)\leq{\mathfrak{c}}$. Then, for all $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$, the ultrapowers $M^{\mathcal{U}}$ and $N^{\mathcal{V}}$ are isomorphic.
@@
By section 3.3, $M^{\mathcal{U}}$ is ${\mathfrak{c}}=\aleph_{1}$-saturated under the Continuum Hypothesis. Thus, by virtue of section 4, it suffices to show that the density character of $M^{\mathcal{U}}$ is equal to $\mathfrak{c}$}. By the construction of the ultraproduct, we have $\chi(M^{\mathcal{U}})\leq\chi(M)^{\aleph_{0}}\leq{\mathfrak{c}}^{\aleph_{0}}={\mathfrak{c}}$. On the other hand, by section 3.1, we can regard $M$ as a substructure of $M^{\mathcal{U}}$, and hence as a closed subset of $M^{\mathcal{U}}$. Thus, if $M$ is non-compact, then $M^{\mathcal{U}}$ is also non-compact. Therefore, by section 3.3, we have that $\chi(M^{\mathcal{U}})\geq\aleph_{1}={\mathfrak{c}}$. Hence, we obtain $\chi(M^{\mathcal{U}})={\mathfrak{c}}$. ∎
Remark 4.4.
In contrast to this result, the following is known: Assume that the Continuum Hypothesis fails. Then, for any infinite-dimensional $\mathrm{C}^{\ast}$}-algebra AA or, more generally, any separable structure that has nonisomorphic ultrapowers associated with ultrafilters in $\beta\omega\setminus\omega$, $A$ has $2^{\mathfrak{c}}$ pairwise nonisomorphic ultrapowers associated with ultrafilters in $\beta\omega\setminus\omega$. See [FHS1:2013] and [FHS2:2014] for details.
@@
Let $M$ and $N$ be saturated models of $\mathrm{T}$ that have the same uncountable density character. Suppose that $A\preceq M$ with $\chi(A)<\chi(M)$ and $\Phi\colon A\to N$ is an elementary embedding. Then, there exist two distinct partial elementary embeddings $\Phi_{0},\Phi_{1}\colon M\to N$ which both extend $\Phi$.
@@
Let $M$ and $N$ be saturated models of $\mathrm{T}$ that have the same uncountable density character. Suppose $A\preceq M$ with $\chi(A)<\chi(M)$, and let $\Phi\colon A\to N$ be an elementary embedding. Then, there exist at least $2^{\chi(M)}$ isomorphisms from $M$ to $N$ that extend $\Phi$.
@@
Assume that the Continuum Hypothesis holds. Suppose $M$ and $N$ are non-compact separable models of $\mathrm{T}$, and let $\Phi\colon M\to N$ be an isomorphism. Then, for any $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$, there exist $2^{\mathfrak{c}}$ isomorphisms from $M^{\mathcal{U}}$ to $N^{\mathcal{V}}$ that extend $\Phi$.
@@
Assume that the Continuum Hypothesis holds. Let $A,B$ be infinite-dimensional separable unital $\mathrm{C}^{\ast}$}-algebras and $\Phi\colon A\to B$ an isomorphism. Then, for any $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$, there exist at least $2^{\mathfrak{c}}$ isomorphisms from $A^{\prime}\cap A^{\mathcal{U}}$ to $B^{\prime}\cap B^{\mathcal{V}}$.
@@ 5 Main Result
In this section, we prove our main result, section 1. We begin by discussing the general setting. Throughout this section, we assume that the language $\mathcal{L}$ is separable with respect to the complete theory $\mathrm{T}$.
Let $X$ be a set equipped with a binary relation $R$. A subset $A\subseteq X$ is said to be cofinal if, for every $x\in X$, there exists $a\in A$ such that $xRa$. The cofinality of $X$, denoted by $\operatorname{cf}(X,R)$, is the least of the cardinalities of the cofinal subsets of $(X,R)$.
Theorem 5.1.
Assume that $\mathrm{T}$ has a definable set $X$ relative to $\mathrm{T}$ and a (possibly definable) formula $\varphi(\bar{x},\bar{y})$ satisfying the following conditions:
  • $\varphi$ takes discrete values $\{0,1\}$ on $X$. That is, for every $M\models\mathrm{T}$, we have $\varphi^{M}(\bar{a},\bar{b})\in\{0,1\}$ for all $\bar{a},\bar{b}\in X(M)$.
  • $\varphi$ defines an asymmetric relation on $X$. That is, for every $M\models\mathrm{T}$ and all $\bar{a},\bar{b}\in X(M)$, if $\varphi^{M}(\bar{a},\bar{b})=0$, then $\varphi^{M}(\bar{b},\bar{a})=1$.
  • For any $M\models\mathrm{T}$, any $n<\omega$, and any $\bar{a}_{i}\in X(M)$ $(i<n)$, there exists $\bar{b}\in X(M)$ such that $\varphi^{M}(\bar{a}_{i},\bar{b})=0$ for all $i<n$.
Then, under the failure of the Continuum Hypothesis, there exist models $A,B\models\mathrm{T}$ with density characters at most $\mathfrak{c}$} such that $A^{\mathcal{U}}\not\cong B^{\mathcal{V}}$ for all $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$.
Proof.
By section 3.3, we take a $\kappa$-saturated model $M$ of $\mathrm{T}$, for large enough $\kappa$. For $\bar{a},\bar{b}\in X(M)$, define $\bar{a}\vartriangleleft\bar{b}$ by $\varphi^{M}(\bar{a},\bar{b})=0$. Let $\bar{a}\trianglelefteq\bar{b}$ denote $\bar{a}\vartriangleleft\bar{b}$ or $\bar{a}=\bar{b}$.
For ordinals $\alpha\leq{\mathfrak{c}}$, we define $\bar{a}_{\alpha}\in M$ and $M_{\alpha}\preceq M$ inductively that satisfy the following conditions:
  • For all $\alpha<{\mathfrak{c}}$, we have $M_{\alpha}\preceq M_{\alpha+1}$ and $\chi(M_{\alpha})\leq\left\lvert{\alpha}\right\rvert+\aleph_{0}$.
  • For all $\alpha<{\mathfrak{c}}$, we have $\bar{a}_{\alpha}\in X(M_{\alpha+1})$ and $\bar{a}\vartriangleleft\bar{a}_{\alpha}$ for all $\bar{a}\in X(M_{\alpha})$.
  • For every limit ordinal $\gamma\leq{\mathfrak{c}}$, we have $M_{\gamma}=\overline{\bigcup_{\alpha<\gamma}M_{\alpha}}$.
One can find a separable elementaly submodel $M_{0}$ of $M$ by section 2.4. Suppose that $M_{\alpha}$ is already defined. Then, by the assumptions on $\varphi$, the set of formulas
$t(\bar{x})\mathrel{:=}\{d(\bar{x},X)\}\cup\left\{\varphi(\bar{a},\bar{x}):\bar{a}\in X(M_{\alpha})\right\}$ (5.1.1)
is a type over $M$. By the saturation of $M$, we can choose $\bar{a}_{\alpha}\in M$ that is the realization of $t(\bar{x})$. By section 2.4, we can take $M_{\alpha+1}\preceq M$ such that $M_{\alpha}\cup\{\bar{a}_{\alpha}\}\subseteq M_{\alpha+1}$ and $\chi(M_{\alpha+1})\leq\chi(M_{\alpha})+\aleph_{0}$.
To prove the statement, it suffices to show the following claim:
Claim.
For every uncountable regular cardinal $\kappa\leq{\mathfrak{c}}$ and every $\mathcal{U}\in\beta\omega\setminus\omega$, we have
$\operatorname{cf}\bigl(X(M_{\kappa}^{\mathcal{U}}),\trianglelefteq\bigr)=\kappa.$ (5.1.2)
(Proof of Claim).
First, we show that $\operatorname{cf}(X(M_{\kappa}^{\mathcal{U}}),\trianglelefteq)\leq\kappa$. Let $C\mathrel{:=}\left\{[(\bar{a}_{\alpha})_{n}]:\alpha<\kappa\right\}\subseteq X(M_{\kappa}^{\mathcal{U}})$. Obviously, $\left\lvert{C}\right\rvert\leq\kappa$. We show that $C$ is cofinal in $(X(M_{\kappa}^{\mathcal{U}}),\trianglelefteq)$. For $\bar{d}\in X(M_{\kappa}^{\mathcal{U}})$, we may assume $\bar{d}=[(\bar{d}_{n})_{n}]$ where $\bar{d}_{n}\in X(M_{\kappa})$ for $n<\omega$. By section 2.4, we can take $\alpha_{n}<\kappa$ such that $\bar{d}_{n}\in X(M_{\alpha_{n}})$ for each $n<\omega$. Let $\alpha\mathrel{:=}\sup_{n<\omega}\alpha_{n}$. Then $\alpha<\kappa$ because $\kappa$ is regular. Put $\bar{c}\mathrel{:=}[(\bar{a}_{\alpha})_{n}]\in C$. One has $\bar{d}_{n}\trianglelefteq\bar{a}_{\alpha}$ for all $n<\omega$. Therefore, by section 3.1, we obtain $\bar{d}\trianglelefteq\bar{c}$.
[…]
Assume that the Continuum Hypothesis fails. Then we have $\omega_{2}\leq{\mathfrak{c}}$. Thus, by the claim above, letting $A\mathrel{:=}M_{\omega_{1}}$ and $B\mathrel{:=}M_{\omega_{2}}$, we obtain
$\operatorname{cf}(X(A^{\mathcal{U}}),\trianglelefteq)=\omega_{1}\neq\omega_{2}=\operatorname{cf}(X(B^{\mathcal{V}}),\trianglelefteq)$ (5.1.3)
for any $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$, thereby completing the proof. ∎
Remark 5.2.
The conditions of section 5 can be described in terms of formulas as follows:
  • $\mathrm{T}\models\sup_{\bar{x},\bar{y}\in X}\min\{\left\lvert{\varphi(\bar{x},\bar{y})}\right\rvert,\left\lvert{1-\varphi(\bar{x},\bar{y})}\right\rvert\}$.
  • $\mathrm{T}\models\sup_{\bar{x},\bar{y}\in X}\max\{0,1-\varphi(\bar{y},\bar{x})-\varphi(\bar{x},\bar{y})\}$.
  • For all $n<\omega$, we have
    $\mathrm{T}\models\sup_{\bar{x}_{0}\in X}\dots\sup_{\bar{x}_{n-1}\in X}\inf_{\bar{y}\in X}\max_{i<n}\varphi(\bar{x}_{i},\bar{y}).$ (5.2.4)
Consequently, since $\mathrm{T}$ is assume to be complete, if these conditions hold in some model of $\mathrm{T}$, then they hold in every model of $\mathrm{T}$.
@@
In order to prove section 1, it suffices to find a theory of unital $\mathrm{C}^{\ast}$}-algebras that satisfies the conditions of section 5. For a $\mathrm{C}^{\ast}$}-algebra AA, $\mathcal{P}(A)$ denotes the set of all projections in $A$. For $\in\mathcal{P}(A)\setminus\{0\}, pp$ is said to be minimal if, for any $q\in\mathcal{P}(A)$ with $0\leq q\leq p$ one has $q=0$ or $q=p$.
@@
Let $\mathrm{T}$ be a theory of unital $\mathrm{C}^{\ast}$}-algebras AA satisfying the following conditions:
  • For all $p,q\in\mathcal{P}(A)$, we have $pq=qp$.
  • AA has no minimal projections. That is, for every $p\in\mathcal{P}(A)\setminus\{0\}$, there exists $q\in\mathcal{P}(A)$ such that $0\lneq q\lneq p$.
Then, there exists a formula satisfies the conditions in section 5.
@@
Let $X(\cdot)\mathrel{:=}\mathcal{P}(\cdot)\setminus\{0,1\}$. We note that $X$ is definable relative to $\mathrm{T}_{{\mathrm{C}^{\ast}}}$, since $X$ is the zero-set of $\left\lVert{x^{*}-x}\right\rVert+\left\lVert{x^{2}-x}\right\rVert+(1-\left\lVert{x}\right\rVert)+(1-\left\lVert{1-x}\right\rVert)$ and it satisfies section 3.4 (3).
Let $\varphi(x,y)$ be the formula defined by
$\varphi(x,y)\mathrel{:=}\max\{1-\left\lVert{xy}\right\rVert,1-\left\lVert{x(1-y)}\right\rVert\}.$ (5.3.5)
Then, by the commutativity of projections, both $pq$ and p(1q)p(1-q) are also projections for all $p,q\in X(A)$. Hence, $\varphi$ takes only the discrete values $\{0,1\}$ on $X$. Moreover, $\varphi$ satisfies the following conditions:
  • For every $n<\omega$ and $p_{i}\in X(A)\ (i<n)$, there exists $q\in X(A)$ such that $\varphi^{A}(p_{i},q)=0$ for all $i<n$.
  • For every $n<\omega$ and $q_{i}\in X(A)\ (i<n)$, there exists $p\in X(A)$ such that $\varphi^{A}(p,q_{i})=1$ for all $i<n$.
Suppose that $n<\omega$ and $p_{i}\in X(A)\ (i<n)$ are given. For each $\sigma\in 2^{n}$, we define
$p_{\sigma}\mathrel{:=}\prod_{i<n}\left[(1-\sigma(i))p_{i}+\sigma(i)(1-p_{i})\right].$ (5.3.6)
If $p_{\sigma}\neq 0$, we can choose $q_{\sigma}\in X(A)$ with $0\lneq q_{\sigma}\lneq p_{\sigma}$. Then, the sum $q\mathrel{:=}\sum_{\sigma}q_{\sigma}$ is in $X(A)$ and satisfies $\varphi^{A}(p_{i},q)=0$ for all $i<n$, because $p_{i}=\sum_{\sigma:\sigma(i)=0}p_{\sigma}$ and $1-p_{i}=\sum_{\sigma:\sigma(i)=1}p_{\sigma}$.
Similarly, suppose that $n<\omega$ and $q_{i}\in X(A)\ (i<n)$ are given. For each $\sigma\in 2^{n}$, we define
$p_{\sigma}\mathrel{:=}\prod_{i<n}\left[(1-\sigma(i))q_{i}+\sigma(i)(1-q_{i})\right].$ (5.3.7)
Since $1=\sum_{\sigma}p_{\sigma}$ there exists $\sigma^{\prime}\in 2^{n}$ such that $p_{\sigma^{\prime}}\in X(A)$. Let $p\mathrel{:=}p_{\sigma^{\prime}}$. Then, we have $\varphi^{A}(p,q_{i})=1$ for all $i<n$.
Now, we define $\psi((x,y),(x^{\prime},y^{\prime}))$ by
$\psi((x,y),(x^{\prime},y^{\prime}))\mathrel{:=}\max\{\varphi(x,y^{\prime}),1-\varphi(x^{\prime},y)\}.$ (5.3.8)
Then $\psi$ also takes values $\{0,1\}$ on $X^{2}$. It is easy to see that $\psi$ is an asymmetric relation on $X^{2}$. By the above properties of $\varphi$, we can see that $\psi$ has the desired conditions. ∎
The following is our main result.
Theorem 5.4.
Suppose that $\mathrm{T}$ satisfies the conditions in section 5. Then, under the failure of the Continuum Hypothesis, there exist unital $\mathrm{C}^{\ast}$}-algebras $A,B\models\mathrm{T}$ with density characters at most $\mathfrak{c}$} such that $A^{\mathcal{U}}\not\cong B^{\mathcal{V}}$ for all $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$.
Proof.
This immediately follows from section 5 and section 5. ∎
@@
Let $X$ be the Cantor space $2^{\omega}$. The $\mathrm{C}^{\ast}$}-algebra $A=C(X)$ is an abelian unital $\mathrm{C}^{\ast}$}-algebra with no minimal projections. Thus, the theory of $A$ satisfies the conditions in section 5. More generally, for any abelian unital $\mathrm{C}^{\ast}$}-algebra $A=C(X)$, where $X$ is a compact Hausdorff space, $A$ has no minimal projections if and only if the Boolean algebra $\operatorname{clop}(X)$ consisting of all clopen subsets of $X$ is atomless.